# Fast Radio Bursts as a probe of gravity on cosmological scales

### v1

@@ Abstract
We explore the potential for improving constraints on gravity by leveraging correlations in the dispersion measure derived from Fast Radio Bursts (FRBs) in combination with cosmic shear. Specifically, we focus on Horndeski gravity, inferring the kinetic braiding and Planck mass run rate from a stage-4 cosmic shear mock survey alongside a survey comprising $10^{4}$ FRBs. For the inference pipeline, we utilise hi_class to predict the linear matter power spectrum in modified gravity scenarios, while non-linear corrections are modelled with HMcode, including feedback mechanisms.
Our findings indicate that FRBs can disentangle degeneracies between baryonic feedback and cosmological parameters, as well as the mass of massive neutrinos. Since these parameters are also degenerate with modified gravity parameters, the inclusion of FRBs can enhance constraints on Horndeski parameters by up to $40$ percent, despite being a less significant measurement. Additionally, we apply our model to current FRB data and use the uncertainty in the $\mathrm{DM}-z$ relation to impose limits on gravity. However, due to the limited sample size of current data, constraints are predominantly influenced by theoretical priors. Despite this, our study demonstrates that FRBs will significantly augment the limited set of cosmological probes available, playing a critical role in providing alternative tests of feedback, cosmology, and gravity. All codes used in this work are made publically available.
@@ 1 Introduction
The Universe entered a phase of accelerated expansion around four billion years ago (Riess et al., 1998). This effect is one of the greatest unsolved mysteries of modern cosmology and understanding the mechanisms behind this dark energy is the proclaimed science goal of the stage-4 cosmological surveys Euclid [^1] https://www.euclid-ec.org/, Rubin-LSST [^2] https://www.lsst.org/ and the Roman telescope [^3] https://roman.gsfc.nasa.gov/. In the standard model of cosmology, dark energy is viewed as a perfect fluid driven by the cosmological constant $\Lambda$. Even though this picture is well motivated at the moment (Tripathi et al., 2017; Planck Collaboration et al., 2020; Escamilla et al., 2024), accurate modelling of the dark energy dynamics is of utmost importance due to its implication of large-scale structure (LSS) formation. The recent DESI findings of a possibly time-dependent dark energy equation of state (DESI Collaboration et al., 2024) further compel us to explore alternative gravity models. The simplest modifications to GR involve a dynamic dark energy fluid driven by an additional scalar degree of freedom in the relativistic field equations. A plethora of these modifications are being actively investigated on the cosmological level, including but not limited to Quintessence (Ratra, 1988), Chameleon (Khoury & Weltman, 2004) and $f(R)$ (Carroll et al., 2004). Such theories, including GR, can be generalised under Horndeski theory (Horndeski, 1974); the most general scalar-tensor theory that has second-order field equations (Kobayashi, 2019).
@@
While WL is a particularly strong probe for the underlying matter distribution and its “clumpiness” specifically (Secco et al., 2022), current and future surveys are limited by uncertainties of baryonic processes (Tröster et al., 2022). This problem will only become more severe with more resolved data and hence lower noise levels unveiling correlations on scales smaller than $1\mathrm{Mpc}$, where baryonic feedback can suppress the power spectrum by up to $60\%$. In particular, the general shape of this suppression is unknown, thus, requiring tight priors in order to be used in cosmological inference. A solution might be the steady rise of Fast Radio Burst (FRB) measurements (Lorimer et al., 2007; CHIME/FRB Collaboration et al., 2021); broad, millisecond transient pulses in the radio frequency range that get dispersed by free electrons along their line of sight and that are consequently highly sensitive to baryons. Even though their origin is still debated (Petroff et al., 2019), the majority of them must be of extragalactic origin due to their high Dispersion Measure (DM). Naturally, they have been proposed as a LSS probe (Zhou et al., 2014a), in particular since the Signal-to-Noise Ratio (SNR) of individual measurements is already ample (Yang & Zhang, 2016). FRBs are versatile; they have been suggested as a tool to find the missing baryon content in the intergalactic medium (Muñoz & Loeb, 2018), or more generally as an $\Omega_{\mathrm{b}}h^{2}$ probe (Walters et al., 2018b). Furthermore, investigating modified gravity with them is not a new concept. Their low-resolution high-SNR nature was utilised in strong gravitational fields before (Adi & Kovetz 2021, Jiang et al. 2024) but the lack of host-identified FRBs prevents a statistically meaningful LSS analysis in the WL regime. The Square Kilometre Array (SKA [^4] https://www.skao.int/, Dewdney et al. 2009) will address this problem at the latest (Hashimoto et al., 2020). Until then, we need to identify the expected constraining power and, in particular, limitations of the currently available modelling tools. So far, investigations focused on how to leverage the averaged cosmological signal, the so-called $\mathrm{DM}-z$ relation (e.g. Zhou et al., 2014b; Walters et al., 2018a; Hagstotz et al., 2022; Macquart et al., 2020; Wu et al., 2022; James et al., 2022; Reischke & Hagstotz, 2023a, b) or the statistics of DM fluctuations (e.g. Masui & Sigurdson, 2015; Shirasaki et al., 2017; Rafiei-Ravandi et al., 2021; Bhattacharya et al., 2021; Takahashi et al., 2021; Reischke et al., 2021, 2022, 2023) to constrain the cosmological standard model.
@@ 2.1 Horndeski theory
As opposed to the standard model, the Lagrangian density in modified gravity depends on the metric tensor $g^{\mu\nu}$ and an additional scalar degree of freedom $\phi$. It can be written as
@@
where $G_{2}$, $G_{3}$, $G_{4}$ and $G_{5}$ are arbitrary functions of $\phi$ and $X$. Bellini & Sawicki (2014) have defined four time-dependent functions $\alpha_{i}$, in addition to the effective Planck mass $M_{\mathrm{Pl}}^{2}$, that fully describe the linear evolution of the background of any sufficiently well-behaved theory [^6] Any theory in which the scalar field has no extreme point in time ($\dot{\phi}\neq 0$). that originates from an arbitrary Lagrangian found in the Horndeski action:
$M_{\mathrm{Pl}}^{2}:=$  $\;2\left(G_{4}+X(G_{5;\phi}-2G_{4;X}-H\dot{\phi}G_{5;X})\right),$
@@
    In Horndeski-type theories, the Planck-Mass is not necessarily a constant. Its evolution through cosmic history is captured by $\alpha_{\mathrm{M}}$. A boundary condition is necessary to well-define the Planck mass. A sensible choice is to use the value it possesses today on small scales $M_{\mathrm{Pl}}^{\mathrm{ini}}$. This is a free parameter if a screening mechanism forces Horndeski theory to reduce to GR on small scales. Otherwise, it must be the observed value $M_{\mathrm{Pl}}^{\mathrm{ini}}=2.176434(24)10^{-8}\,\mathrm{kg}$ (Tiesinga et al., 2021). $\alpha_{\mathrm{M}}\neq 0$ indicates a non-minimally coupled theory of gravity [^7] Meaning the “charge” of gravity (the metric $g_{\mu\nu}$) has a more complicated relationship to the scalar field $\phi$ than linear scaling.. In that case, $\alpha_{\mathrm{M}}$ creates anisotropic stress.
@@
    This term measures the mixing between the canonical kinetic terms of the scalar field and the metric. It is particularly interesting for the scope of this thesis because it modifies the growth of perturbations and the shape of the power spectrum, by causing “clustering” of dark energy (Bellini & Sawicki, 2014).
@@
    $\left|1-c_{\mathrm{grav}}/c_{\mathrm{light}}\right|<3\times 10^{-15}\;.$ (5)
    Therefore, the tensor speed excess is largely constrained to its GR value: $\alpha_{\mathrm{T}}=0$.
@@
Well aware that this parameterisation can break down at redshifts $z>10$ (Linder, 2017), a strong motivation for Horndeski theory is whether the late time accelerated expansion of the Universe can (partly) be explained with a non-GR gravity on large cosmic scales. The energy density is driven by the same functions as the $\alpha_{i}$, so we could naively expect $\hat{\alpha}_{i}\sim\mathcal{O}(1)$ (Bellini et al., 2016). Thus, since the kineticity affects LSS evolution only marginally at best, we adapt henceforth $\hat{\alpha}_{\mathrm{K}}=1$. Considering the already strongly constrained $\alpha_{\mathrm{T}}$, the Horndeski theory cosmological parameters inferred in this work will be the braiding and Planck mass run rate proportionality constants, $\hat{\alpha}_{\mathrm{B}}$ and $\hat{\alpha}_{\mathrm{M}}$, respectively. Note that we do not vary the dark energy equation of state here explicitly since our focus is on the modifications to GR on the perturbation level and not at the background.
@@
with the Fourier transformed matter density $\tilde{\rho}_{\mathrm{m}}(\boldsymbol{k},a)$ and wave-number $k:=|\boldsymbol{k}|$. Note that for ${\mu},{\eta}\rightarrow 1$, GR is retrieved. We know that GR is an excellent description of gravity for very small scales at the very least (Wojtak et al., 2011), hence we force modified gravity to reduce to GR for scales greater than a screening scale $k_{\mathrm{s}}$ that we implement as described in Reischke et al. (2019).
@@ 2.2 Angular power spectrum
Figure 1: Tomographic source redshift distribution of the mock FRB catalogue.
@@ 2.2.1 Weak lensing statistics
See Spurio Mancini et al. (2019) for a more detailed derivation, but be aware of the different naming conventions of the Baradeen potentials. Inserting Eq. 15 with $k=\frac{\ell+1}{2}$ into Eq. 13 with the matter power spectrum $P_{\mathrm{mm}}$ yields the auto-correlated cosmic shear angular power spectrum matrix in modified gravity.
The noise component of the observed shear spectra, Eq. (11), called shape noise is given as
@@
2.2.2 Dispersion measure statistics
$\mathrm{DM}=\int_{0}^{L}\mathrm{d}l\,n_{\mathrm{e}}(l)\;.$ (17)
In this work, we investigate DM fluctuations $\mathcal{D}(\hat{n},z)$ which enables us to employ a zero-mean Gaussian field assumption. The corresponding auto-correlated angular power spectrum is then retrieved from Eq. (13) with the electron power spectrum $P_{\mathrm{ee}}$ and the DM kernel
with
$\mathcal{A}=\frac{3H_{0}^{2}\Omega_{\mathrm{b}0}}{8\pi Gm_{\mathrm{p}}}\;,$ (21)
However, for the considered redshift range in this work ($z\leq 4$), the intergalactic medium can be assumed to be fully ionised (Aghanim, 2001). Additionally, the majority of the DM signal comes from $20.5<z<$ 2 (see also the FRB mock source distribution in Section 3.2.1) at which we assume a constant $f_{\mathrm{IGM}}\approx 0.9$. The exact distribution of baryons is still an object of research (Shull et al., 2012) and a more complex model can be implemented in future work. Combining this assumption with the vanishing fraction of heavy elements ($Y_{\mathrm{H}}\approx 1-Y_{\mathrm{He}}$), Eq(22) reduces to
A proper model of the MW and host galaxy impact as noise terms is equally important. Extensive work has been conducted on the MW contribution (Cordes & Lazio, 2002; Yao et al., 2017; Yamasaki & Totani, 2020) and we assume it can be accurately modelled and subtracted. Meanwhile, the host contribution acts as a Poisson-like flat noise contribution that can be characterised by the host DM dispersion
The factor $1+z$ originates from the transformation between the FRB rest- and observer frame (Zhang et al., 2020) and we adapt the host dispersion from Arcus et al. (2021). Thus, we can write down the noise contribution as
$P_{\mathrm{em}}(k,z)=\sqrt{P_{\mathrm{ee}}(k,z)P_{\mathrm{mm}}(k,z)}$ (27)
in addition to WL (Eq. 15) and DM (Eq. 20) kernel into Eq. 13.

### v2

@@ Abstract
We explore the potential for improving constraints on gravity by leveraging correlations in the dispersion measure derived from Fast Radio Bursts (FRBs) in combination with cosmic shear. Specifically, we focus on Horndeski gravity, inferring the kinetic braiding and Planck mass run rate from a stage-4 cosmic shear mock survey alongside a survey comprising $10^{4}$ FRBs. For the inference pipeline, we utilise the Boltzmann code hi_class to predict the linear matter power spectrum in modified gravity scenarios, while non-linear corrections are obtained from the halo-model employed in HMcode, including feedback mechanisms. Our findings indicate that FRBs can disentangle degeneracies between baryonic feedback and cosmological parameters, as well as the mass of massive neutrinos. Since these parameters are also degenerate with modified gravity parameters, the inclusion of FRBs can enhance constraints on Horndeski parameters by up to $40$ percent, despite being a less significant measurement. Additionally, we apply our model to current FRB data and use the uncertainty in the $\mathrm{DM}-z$ relation to impose limits on gravity. However, due to the limited sample size of current data, constraints are predominantly influenced by theoretical priors. Despite this, our study demonstrates that FRBs will significantly augment the limited set of cosmological probes available, playing a critical role in providing alternative tests of feedback, cosmology, and gravity. All codes used in this work are made publically available.
@@ 1 Introduction
The Universe entered a phase of accelerated expansion around four billion years ago (Riess et al., 1998). This effect is one of the greatest unsolved mysteries of modern cosmology and understanding the mechanisms behind this dark energy is the proclaimed science goal of the stage-4 cosmological surveys Euclid [^1] https://www.euclid-ec.org/, Rubin-LSST [^2] https://www.lsst.org/ and the Roman telescope [^3] https://roman.gsfc.nasa.gov/. In the standard model of cosmology, dark energy is viewed as a perfect fluid driven by the cosmological constant $\Lambda$. Even though this picture is well motivated at the moment (Tripathi et al., 2017; Planck Collaboration et al., 2020b; Escamilla et al., 2024), accurate modelling of the dark energy dynamics is of utmost importance due to its implication of large-scale structure (LSS) formation. The recent DESI findings of a possibly time-dependent dark energy equation of state (DESI Collaboration et al., 2025) further compel us to explore alternative gravity models. The simplest modifications to GR involve a dynamic dark energy fluid driven by an additional scalar degree of freedom in the relativistic field equations. A plethora of these modifications are being actively investigated on the cosmological level, including but not limited to Quintessence (Ratra, 1988), Chameleon (Khoury & Weltman, 2004a) and $f(R)$ (Carroll et al., 2004). Such theories, including GR, can be generalised under Horndeski theory (Horndeski, 1974); the most general scalar-tensor theory that has second-order field equations (Kobayashi, 2019).
@@
While WL is a particularly strong probe for the underlying matter distribution and its “clumpiness” specifically (Secco et al., 2022), current and future surveys are limited by uncertainties of baryonic processes (Tröster et al., 2022). This problem will only become more severe with more resolved data and hence lower noise levels unveiling correlations on scales smaller than $1\mathrm{Mpc}$, where baryonic feedback can suppress the power spectrum by up to $40\%$(Chisari et al., 2019a). In particular, the general shape of this suppression is unknown, thus, requiring tight priors in order to be used in cosmological inference. A solution might be the steady rise of Fast Radio Burst (FRB) measurements (Lorimer et al., 2007; CHIME/FRB Collaboration et al., 2021); broad, millisecond transient pulses in the radio frequency range that get dispersed by free electrons along their line of sight and that are consequently highly sensitive to baryons. Even though their origin is still debated (Petroff et al., 2019), the majority of them must be of extragalactic origin due to their high Dispersion Measure (DM). Naturally, they have been proposed as a LSS probe (Zhou et al., 2014), in particular since the Signal-to-Noise Ratio (SNR) of individual measurements is already ample (Yang & Zhang, 2016). FRBs are versatile; they have been suggested as a tool to find the missing baryon content in the intergalactic medium (Muñoz & Loeb, 2018), or more generally as an $\Omega_{\mathrm{b}}h^{2}$ probe (Walters et al., 2018). Furthermore, investigating modified gravity with them is not a new concept. Their low-resolution high-SNR nature was utilised in strong gravitational fields before (Adi & Kovetz 2021, Jiang et al. 2024) but the lack of host-identified FRBs prevents a statistically meaningful LSS analysis in the WL regime. The Square Kilometre Array (SKA [^4] https://www.skao.int/, Dewdney et al. 2009) will address this problem at the latest (Hashimoto et al., 2020). Until then, we need to identify the expected constraining power and, in particular, limitations of the currently available modelling tools. So far, investigations focused on how to leverage the averaged cosmological signal, the so-called $\mathrm{DM}-z$ relation (e.g. Zhou et al., 2014; Walters et al., 2018; Hagstotz et al., 2022; Macquart et al., 2020; Wu et al., 2022; James et al., 2022; Reischke & Hagstotz, 2023a, b) or the statistics of DM fluctuations (e.g. Masui & Sigurdson, 2015; Shirasaki et al., 2017; Rafiei-Ravandi et al., 2021; Bhattacharya et al., 2021; Takahashi et al., 2021; Reischke et al., 2021, 2022, 2023) to constrain the cosmological standard model.
@@ 2.1 Horndeski theory
As opposed to the standard cosmological model, the Lagrangian density in Horndeski modified gravity depends on the metric tensor $g^{\mu\nu}$ and an additional scalar degree of freedom $\phi$(Horndeski, 1974). It can be written as
@@
where $G_{2}$, $G_{3}$, $G_{4}$ and $G_{5}$ are arbitrary functions of $\phi$ and $X$. Bellini & Sawicki (2014) have redefined the $G_{i}$ into four time-dependent functions $\alpha_{i}$ that fully describe linear perturbations to the Horndeski action shown in Eq. 2. With the help of the now time dependent Planck mass
[…]
they can be written as:
@@
    In Horndeski-type theories, the Planck-Mass is not necessarily a constant. Its evolution through cosmic history is captured by $\alpha_{\mathrm{M}}$. A boundary condition is necessary to well-define the Planck mass. A sensible choice is to use the value it possesses today on small scales $M_{\mathrm{Pl}}^{\mathrm{ini}}$. This is a free parameter if a screening mechanism forces Horndeski theory to reduce to GR on small scales. Otherwise, it must be the observed value $M_{\mathrm{Pl}}^{\mathrm{ini}}=2.176434(24)\cdot 10^{-8}\,\mathrm{kg}$ (Tiesinga et al., 2021). $\alpha_{\mathrm{M}}\neq 0$ indicates a non-minimally coupled theory of gravity [^6] Meaning the “charge” of gravity (the metric $g_{\mu\nu}$) has a more complicated relationship to the scalar field $\phi$ than linear scaling.. In that case, $\alpha_{\mathrm{M}}$ creates anisotropic stress.
@@
    This term measures the mixing between the canonical kinetic terms of the scalar field and the metric. It is particularly interesting for the scope of this paper because it modifies the growth of perturbations and the shape of the power spectrum, by causing “clustering” of dark energy (Bellini & Sawicki, 2014).
@@
    $\left|1-c_{\mathrm{grav}}/c_{\mathrm{light}}\right|<3\times 10^{-15}$ (5)
    on scales LIGO (LIGO Scientific Collaboration et al., 2015) is sensitive to. However, the tensor speed excess could be in principle scale dependent due to various screening mechanisms and the implications of this measurement as still debated (De Rham & Melville, 2018; Noller, 2020). The Laser Interferometer Space Antenna (LISA, Auclair et al. 2023) will likely be sensitive enough to give conclusive measurements of $\alpha_{\text{T}}$ on cosmic scales. Nevertheless, we employ the assumption here that $\alpha_{\mathrm{T}}=0$ for linear perturbations. This rules out a substantial amount of modified gravity theories acting on dark energy. Still, popular theories like quintessence (Ratra & Peebles, 1988) and $f(R)$(Carroll et al., 2004) remain viable (Ezquiaga & Zumalacárregui, 2017).
@@
Well aware that this parameterisation can break down at redshifts $z>10$ (Linder, 2017), a strong motivation for Horndeski theory is whether the late time accelerated expansion of the Universe can (partly) be explained with a non-GR gravity on large cosmic scales. The energy density is driven by the same functions as the $\alpha_{i}$, so we could naively expect $\hat{\alpha}_{i}\sim\mathcal{O}(1)$ (Bellini et al., 2016). Thus, since the kineticity affects LSS evolution only marginally at best, we adapt henceforth $\hat{\alpha}_{\mathrm{K}}=1$. Considering the constraints on $\alpha_{\mathrm{T}}$, the Horndeski theory cosmological parameters inferred in this work will be the braiding and Planck mass run rate proportionality constants, $\hat{\alpha}_{\mathrm{B}}$ and $\hat{\alpha}_{\mathrm{M}}$, respectively. Note that we do not vary the dark energy equation of state here explicitly since our focus is on the modifications to GR on the perturbation level and not at the background.
@@
Figure 1: Tomographic source redshift distribution of the mock FRB catalogue. The discontinuities between bins originate from normalising each source distribution individually.
with the Fourier transformed matter density $\tilde{\rho}_{\mathrm{m}}(\boldsymbol{k},a)$ and wave-number $k:=|\boldsymbol{k}|$. Note that for ${\mu},{\eta}\rightarrow 1$, GR is retrieved. We know that GR is an excellent description of gravity for very small scales at the very least (Wojtak et al., 2011), hence we force modified gravity to satisfy GR on these scales by the means of screening. There are a variety of proposed screening mechanisms (see Joyce et al. 2015 for an excellent review thereof). The ones most commonly discussed in literature are Chameleon (Khoury & Weltman, 2004a, b) and Vainshtein screening (Vainshtein, 1972; Babichev & Deffayet, 2013). In Chameleon-like screening, the mass of the scalar field depends on the local matter density, making the effects of modified gravity for high-density environments like Earth short-ranged, hence the name. On the other hand, Vainshtein screening suppresses scalar field effects in the presence of massive sources where the $\phi$-kinetic terms becomes large. This happens on scales smaller than a characteristic distance named the Vainshtein radius, which is dependent on the properties of the gravity source. In this work, we employ a phenomenological model described in Reischke et al. (2019); Spurio Mancini et al. (2019), where $\eta(k,a)$ and $\mu(k,a)$ decay exponentially to unity for scales smaller than the screening scale $\lambda_{\mathrm{s}}\sim 1/k_{\mathrm{s}}$. In this work, we refrain from testing the effects of implementing a variety of screening mechanisms. However, preliminary results suggest that the approach taken here does not influence the power spectrum up to a few per cent (Grasso et al., in prep.).
One would expect the screening scale to affect extremely non-linear scales and/or high power spectrum amplitudes with the Chameleon or Vainshtein screening. However, we restrict the screening scale to be below intermediate scales, $k_{\mathrm{s}}<2\,\mathrm{Mpc}^{-1}$ with a fiducial value at $k_{\mathrm{s}}=0.1\,\mathrm{Mpc}^{-1}$. The reason for only considering (semi-)linear scales is of practical nature: We use HMcode to model the non-linear power spectrum (see also Section 3.3). Notably, HMcode was developed and calibrated for a $\Lambda$CDM cosmology and has no inherent knowledge and functionality of modified gravity. While there are efforts to model the non-linear matter power spectrum in scalar-tensor perturbation theory (Cataneo et al., 2019; Bose et al., 2023), there is little similar quantitative research for the baryon power spectrum. To circumvent this issue we remove modified gravity for large $k$, avoiding the necessity of modelling the effect of modified gravity on baryons. Therefore, including modified gravity on non-linear scales would only increase information gain therein, leaving the results of this work valid as a lower bound.
@@ 2.2.1 Weak lensing statistics
See Spurio Mancini et al. (2019) for a more detailed derivation, but be aware of the different naming conventions of the Bardeen potentials. The quasi-static approximation (QSA) is often used in the perturbative Einstein equations, where time derivatives are neglected, allowing for an explicit relationship between the $\alpha_{i}$ and $\mu(k,\chi)$ as well as $\eta(k,\chi)$(Sawicki & Bellini, 2015). However, the hi_class modification we use (see also Ch. 3.3) does not employ the QSA. Instead, the density and potential are sourced directly from hi_class that considers the full statistical properties of the Weyl potential (Spurio Mancini et al., 2018b). Inserting Eq. 15 with $k=\frac{\ell+1}{2}$ into Eq. 13 with the matter power spectrum $P_{\mathrm{mm}}$ yields the auto-correlated cosmic shear angular power spectrum matrix in modified gravity.
The noise component of the observed shear spectra, Eq. 11, called shape noise is given as
@@
2.2.2 Dispersion measure statistics
Parameter description fiducial value prior range source
$\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ Horndeski theory screening scale -1 [$-2$, $0.3$] Spurio Mancini et al. (2018b)
$h$ dimensionless Hubble parameter 0.674 [$0.38$, $1.00$] Planck Collaboration et al. (2020b)
$\Omega_{\mathrm{b}}$ baryon density parameter $0.04931$ [$0.015$, $0.1$] Planck Collaboration et al. (2020b)
$\Omega_{\mathrm{cdm}}$ cold dark matter density parameter $0.2642$ [$0.18$, $0.34$] Planck Collaboration et al. (2020b)
$n_{\mathrm{s}}$ primordial scalar index $0.965$ [$0.7$, $1.25$] Planck Collaboration et al. (2020b)
$\sum m_{\nu}/\mathrm{eV}$ total mass of all neutrino species $0.06$ [$0.003$, $1.5$] Planck Collaboration et al. (2020b)
$\sigma_{8}$ linear power spectrum RMS fluctuations $0.811$ [$0.7$, $0.92$] Planck Collaboration et al. (2020b)
fixed values $N_{\nu}$ number of neutrino species 3 Planck Collaboration et al. (2020b)
$M_{\mathrm{Pl}}$ initial Planck mass $M_{\mathrm{Pl}}^{\mathrm{obs}}$
$\hat{\alpha}_{\mathrm{K}}$ kineticity function proportionality $1$
$N_{\mathrm{eff}}$ effective relativistic degrees of freedom 3.046 Planck Collaboration et al. (2020b)
$T_{0}$ Black body CMB temperature today $2.7255\,\mathrm{K}$ Fixsen (2009)
$z_{\mathrm{reio}}$ reionisation redshift 7.7 Planck Collaboration et al. (2020b)
$Y_{\mathrm{He}}$ primordial helium fraction 0.246 Planck Collaboration et al. (2020b)
$w(a)$ dark energy equation of state -1 Planck Collaboration et al. (2020b)
$\Omega_{\mathrm{K}}$ curvature density parameter 0 Planck Collaboration et al. (2020b)
$\mathrm{DM}=\int_{0}^{L}\mathrm{d}l\,n_{\mathrm{e}}(l)\;.$ (17)
In this work, we investigate DM fluctuations $\mathcal{D}(\hat{n},z)$ which enables us to employ a zero-mean Gaussian field assumption. The corresponding auto-correlated angular power spectrum is then retrieved from Eq. 13 with the electron power spectrum $P_{\mathrm{ee}}$ and the DM kernel
with
$\mathcal{A}=\frac{3H_{0}^{2}\Omega_{\mathrm{b}0}}{8\pi Gm_{\mathrm{p}}}\;,$ (21)
However, for the considered redshift range in this work ($z\leq 4$), the intergalactic medium can be assumed to be fully ionised (Aghanim, 2001). Additionally, the majority of the DM signal comes from $20.5<z<$ 2 (see also the FRB mock source distribution in Section 3.2.1) at which we assume a constant $f_{\mathrm{IGM}}\approx 0.9$. The exact distribution of baryons is still an object of research (Shull et al., 2012) and a more complex model can be implemented in future work. Combining this assumption with the vanishing fraction of heavy elements ($Y_{\mathrm{H}}\approx 1-Y_{\mathrm{He}}$), Eq. 22 reduces to
A proper model of the MW and host galaxy impact in terms of noise is equally important. Extensive work has been done on the contribution of the MW (Cordes & Lazio, 2002; Yao et al., 2017; Yamasaki & Totani, 2020) and we assume here that it can be accurately modelled and subtracted. In reality this might not be the case and improper $\text{DM}_{\text{MW}}$ modelling will lead to large-scale contributions of the DM-DM auto-correlation. However, since shear and $\text{DM}_{\text{MW}}$ are uncorrelated, we expect the gain in constraining power that is driven by the cross-correlation largely unaffected. Furthermore, selecting FRBs outside of the MW disk would reduce the possible residual non-LSS DM contributions. Meanwhile, the host contribution acts as a Poisson-like flat noise contribution that can be characterised by the host DM dispersion
The factor $1+z$ originates from the transformation between the FRB rest- and observer frame (Zhang et al., 2020). The host dispersion is still a large source of uncertainty (Tang et al., 2023). Simulations suggest that the $\sigma_{\text{host0}}$ is rather on order of $\sim$100 pc cm-3(Kovacs et al., 2024; Zhang et al., 2024). Still, here we fix the contribution to the observationally motivated value from Arcus et al. (2021). Even though the assumed host dispersion measure might be optimistic, it is highly degenerate with the total number of FRBs (Reischke et al., 2023, Fig. 4 therein). Recent estimates by Connor & Ravi (2023) suggest that the supposed figure of 10000 FRBs until the end of the decade might be on the low end. Thus, we do not expect our choice of parameters to affect the conclusion of this paper. A more rigorous approach would be to additionally marginalise over the host DM dispersion with an analytical model, like we propose in Reischke et al. (2024). However, this is outside the scope of this paper, hence we choose a simplistic host dispersion model. As such, we can write down the noise contribution as
$P_{\mathrm{em}}(k,z)=\sqrt{P_{\mathrm{ee}}(k,z)P_{\mathrm{mm}}(k,z)}$ (27)
in addition to WL (Eq. 15) and DM (Eq. 20) kernel into Eq. 13.